# Position: A taxonomy for reporting and describing AI security incidents

### v1

@@ Abstract
AI systems are vulnerable to attacks, and corresponding AI security incidents have been described. Although a collection of safety incidents around AI will become a regulatory requirement, there is no proposal to collect AI security incidents. In this position paper, we argue that a proposal should be made, taking into account the interests and needs of different stakeholders: industry, providers, users, and researchers. We thus attempt to close this gap and propose a taxonomy alongside its requirements like machine readability and link-ability with existing databases. We aim to spark discussions and enable discussion of which information is feasible, necessary, and possible to report and share within and outside organizations using AI.
@@ 1 Introduction
Work in progress, shared for feedback purposes only. Authorlist non-final.
@@
Artificial Intelligence (AI) is vulnerable to security threats (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). Recently, evidence appeared that such attacks occur on real-world systems (Grosse et al., 2024b; McGregor, 2021; ARVA, ). Such reports include, but are not limited to, malicious AI dependencies [^1] https://avidml.org/database/, 2023-V015, bypassed malware detectors [^2] https://avidml.org/database/, 2023-V004, failing perception modules of autonomous vehicles and reverse engineering of existing models (Grosse et al., 2024b), and search engine poisoning [^3] https://incidentdatabase.ai/cite/88.
Reporting and analyzing any incidents related to AI systems is a fundamental prerequisite for adequate risk management. For non-security incidents in context of AI systems, there are existing efforts for monitoring (OECD, ; Huntr, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, b, a; Walker et al., 2024). Such databases enable analysis for deeper understanding (Kumar & Sharma, 2022). While there are taxonomies that support the reporting of cybersecurity vulnerabilities in software more broadly, such as the Common Vulnerability and Exposure (CVE) program (MITRE, e; Baker et al., 1999), these taxonomies do not apply well to AI. In contrast to shareable and protected software, AI model specifics (such as architecture, learned weights, and hyperparameters) may affect vulnerabilities (Biggio & Roli, 2018; Gu et al., 2019), but may at the same time be the intellectual property or business asset of a company (Grosse et al., 2023). Additionally, the CVE program is monitors vulnerabilities. In contrast, AI model vulnerabilities are often well-known but the details of their practical exploitation are not. AI security incident reporting requires more context for understanding AI-specific vulnerabilities and exploits that are not covered by conventional incident reporting.
Meanwhile, AI incident reporting is a developing obligation according to several regulatory frameworks. The European AI Act, as exemplary horizontal legislation, defines ’serious incidents’ in context of AI systems and requires reporting to the European AI Office and national competent authorities. At the same time, there are no corresponding templates and taxonomies for AI security incident reporting. A consensus on the necessary content, scope and granularity of suchlike frameworks has not been reached from a scientific, industry and market surveillance perspective.
@@
Our position: We argue that a reporting taxonomy for AI security incidents is necessary and (1) should contain specific information on the AI system, the incident and its implications. (2) Its structure should be aligned with existing reporting approaches and existing AI (security) governance. (3) The content should be of sufficient granularity for industrial practitioners, researchers and market surveillance authorities to derive required action or, in some cases, even reproduce the incident. The taxonomy should further be usable, and machine readable, properties that however depend on the exact implementation. We present this draft to be further discussed by the community.
Outline. We define all relevant terminology based on existing standards or works in Section 2.1. Based on these definitions, we review generally related concepts and approaches in Section 2. We then describe a potential taxonomy for AI security incident reporting in Section 3: We first collected feedback from twelve experts for a draft reporting scheme, which was then iteratively improved. Afterwards, we describe the resulting scheme in Section 4. We then discuss limitations and implications in Section 5. For example, while our taxonomy does not cover offensive AI or recursive AI usage, we outline important implementation details. These include desired properties, such as machine readability, but also reproducibility, and the use of existing frameworks for attacks and harm classifications. To conclude, we identify open research questions like the usability of AI security reporting frameworks to ease application, and the question how to extend the framework to potentially collect information on the attacker. Before we conclude with the impact statement, we discuss alternative views in Section 6.
@@
Related Work. To the best of our knowledge, there is no established procedure for AI security incident reporting. The closest approach to our work is by Cattell & Ghosh (2024). They propose a CVE-like framework for adjudicating AI flaws more generally (e.g., incident reporting for undesired behaviors of AI systems which is not focused on security) as well as other improvements (such as extended model cards). In this sense, Cattell & Ghosh (2024) is orthogonal to our work. In addition, MITRE announced a general AI incident sharing initiative [^4] https://ai-incidents.mitre.org concurrently to our work. While their proposal bears similarities to ours, we suggest a more detailed treatment of the intended purpose, supply chains, and existing security measures.
@@ 2 Background
In this section, we first introduce relevant definitions before we review existing works in the area of describing AI systems, AI security, as well as reporting AI and security incidents.
@@ 2.1 Definitions
As AI incident reporting is also increasingly related to regulatory requirements, we seek to define all involved terms based on existing definitions where possible. The complete list of all required definitions is in Appendix A. Here, we only revisit the three most important definitions: firstly an “AI system” (the one the incident is reported of), secondly the “CIA” security triad, and thirdly an “AI security incident” (the target of our collection). We first start with the definition of the AI system, derived from the OECD (2024a) definition.
@@
AI system. An AI model embedded into a larger system, together with its supporting and surrounding software and hardware. An AI system can influence physical or virtual environments. Different AI systems vary in their levels of autonomy and adaptiveness after deployment.
The incident collection we propose is limited to systems falling under this definition and covers only AI security incidents. We thus also define security as defined by ISO (2018).
@@
We further additionally review the definition of confidentiality, integrity, and availability, better known as the CIA triad, and also well established in scientific work covering AI security (Biggio & Roli, 2018).
@@
Based on these security terms, we define AI security based on existing definitions by the OECD (2024b) and ISO [^5] ISO/IEC TR 24028 Information technology — Artificial intelligence — Overview of trustworthiness in artificial intelligence.
@@
Our proposed taxonomy to collect AI security incidents is only intended for the above-encompassed incidents. We use these definitions to avoid, for example, confusion between security and AI security incidents (Bieringer et al., 2022). Also note that the current definition above entails potentially affected, hence vulnerabilities or red teaming results. It may be desirable to formulate this more strictly later-on.
@@ 2.2 Describing AI systems
Since the reporting of security incidents should be consistent with the overall documentation of an AI system, our proposed taxonomy builds on and expands prior work on model cards. After their initial proposal by Mitchell et al. (2019), these have been widely adopted by model providers and are used by various stakeholder groups to access information about a model and conclude corresponding action. Derivative templates have been proposed as fact sheets (Arnold et al. (2019)), datasheets for datasets (Gebru et al. (2021)), method cards (Adkins et al. (2022)), risk cards (Derczynski et al. (2023)), use case cards (Hupont et al. (2024)), and AI cards (Golpayegani et al. (2024)). While most of these approaches require the documentation of general metadata about the model (e.g. model type, model version), their content related to security is oftentimes limited and only covers implemented security measures, if at all.
@@ 2.3 AI security
Considering security is however crucial, as similar to any other (digital) system, an AI system targeted by a plethora of attacks (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). In the following, we give a very high-level overview, based on the previously discussed CIA triad, of attacks later used to sanity check the proposed reporting framework.
Poisoning attacks alter the training data, samples, or labels to decrease the overall performance of the classifier (Cinà et al., 2023; Zhang et al., 2020). They, thus, target the integrity via the performance, or availability of the AI model if performance becomes too low.
Table 1: Existing databases and collections of AI (security) incidents. We list the kind of support party (Government, Industry, Non-Profit, and Academia) and incident numbers as of June 14th, 2024. We list the number of entries in each database and whether incidents cover the topics of ethics, security, AI security, or offensive use of AI within security. When then review which databases contain reproducible incidents. We then revisit the sources of the incidents: research papers, media, anonymous reports, or platforms where registered users can submit reports. We finally report whether the databases are machine-readable and linked with the CVE databases. The AVID database is linked to NVD, MITRE, and AIID.
@@
MITRE case studies (MITRE, b) G 22 JSON Python CVE
AIRD (MITRE, a) I 4 CSV Python CVE
Model stealing copies the model without the owner’s consent (Oliynyk et al., 2023; Chen et al., 2021; Carlini et al., ), thus targeting confidentiality.
@@
Many more attacks, including backdoors (Cinà et al., 2023; Yu et al., 2022; Yang et al., 2024), weight perturbations (Neti et al., 1992; Breier et al., 2018), adversarial initialization (Grosse et al., 2020; Liu et al., 2019), and sponge attacks (Shumailov et al., 2020) have been discussed. The exact nature and scope of describing and tagging a specific attack in a practical context is a research question itself and thus left for future work. There are, however, established frameworks like the MITRE ATLAS [^6] https://atlas.mitre.org/matrices/ATLAS addressing this challenge in practice.
2.4 AI incidents
@@
Although there are rather few reports of real-world AI security incidents as discussed in the introduction (Grosse et al., 2024b; McGregor, 2021; ARVA, ), there are many reports of general incidents in the context of AI systems. We summarize related databases in Table 1. Most databases cover not only AI security incidents but instead a range of incidents, ranging from unethical applications of AI to security, and offensive use of AI (OECD, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, b). Very few databases are more specific and focus on non-AI security (Huntr, ) or offensive AI usage (Walker et al., 2024). In addition, most databases rely on publicly available information, like research papers and media reports (OECD, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, b). Far fewer rely (also) on direct submissions (Huntr, ; LVEs, ; MITRE, a) in the spirit of bug bounties. Even fewer use anonymous reports (Grosse et al., 2024b). Each format constrains the kind of information that is available per database entry. While some databases, such as the AI Incident Database (AIID) (McGregor, 2021), apply taxonomies to incidents, they do not require the information needed for an in-depth security analysis that allows to determine if other AI systems are vulnerable or not. Opposed to this, MITRE recently released an AI incident sharing initiative [^7] https://ai-incidents.mitre.org collecting more information.
@@
Still, these databases provide guidelines on collected information and emphasize other relevant features. These include machine readability (seven of nine databases allow direct export of data), links to other vulnerability databases like CVE (four of nine), and lastly direct integration via for example APIs with programming languages for data analysis (five of nine). The latter is especially relevant to research, where corresponding frameworks in non-AI security have enabled more detailed analysis of the threat surface (Neuhaus & Zimmermann, 2010; Li et al., 2017).
2.5 Existing infrastructure for security incidents
Within security, frameworks to collect vulnerabilities exists. The main reporting infrastructure for security is the CVE program (Baker et al., 1999). Similar taxonomies have evolved over time to accommodate emerging new methods and tools (MITRE, f). For example, in the case of CVE (Baker et al., 1999), a related method, the Common Vulnerability Scoring System (CVSS) (NIST, a; Scarfone & Mell, 2009), was developed to assess the severity of vulnerabilities. As a result, the National Vulnerability Database (NVD) (NIST, b) now contains CVSS scores for all published CVE records. Additionally, alongside the descriptions of vulnerabilities and exposures, the need to enumerate attacks was identified. The MITRE ATT&CK framework (MITRE, c, d), which serves as a taxonomy of attack tactics and techniques, was created to address this gap.
Previous works outline that security and AI security vulnerabilities are fundamentally different (NIST, 2023; Cattell & Ghosh, 2024). We agree with this position, yet acknowledge that CVEs can be relevant for AI models [^8] For an example, consider CVE-2023-25661.. CVEs are however orthogonal to our reporting in that they cover model independent vulnerabilities in for example code, while our taxonomy covers model specific vulnerabilities (which may extend to several models, but are not fully model agnostic).

### v2

@@ Abstract
As AI usage becomes more ubiquitous, AI incident reporting is both practiced increasingly in industry and mandated by regulatory requirements. At the same time, it is established that AI systems are exploited in practice by a growing number of security threats. Yet, organizations and practitioners lack necessary guidance in describing AI security incidents. In this position paper, we argue that specific taxonomies are required to describe and report security incidents of AI systems. In other words, existing frameworks for either non-AI security or generic AI safety incident reporting are insufficient to capture the specific properties of AI security. To demonstrate our position, we offer an AI security incident taxonomy and highlight relevant properties, such as machine readability and integration with existing frameworks. We have derived this proposal from interviews with experts, aiming for standardized reporting of AI security incidents, which meets the requirements of affected stakeholder groups. We hope that this taxonomy sparks discussions and eventually allows the sharing of AI security incidents across organizations, enabling more secure AI.
@@ 1 Introduction
Reporting and analyzing any incidents related to Artificial Intelligence (AI) systems is a fundamental prerequisite for adequate risk management. For non-security incidents in the context of AI systems, there are existing efforts for monitoring incidents (OECD, b; Huntr, ; McGregor, 2021; ARVA, ; LVEs, ; MITRE, a, ; Walker et al., 2024). In other areas like aviation, the rigorous study of incidents has lead to a decrease in casualties over time (Li, 2019). Consequently, AI incident reporting is becoming a relevant obligation according to regulatory frameworks. The European AI Act, as exemplary horizontal legislation, requires reporting of serious incidents to the European AI Office and national competent authorities. Yet, this entails a broad range of incidents: general incidents as above, but also incidents where an attacker manipulates an AI system.
@@
There is a plethora of scientific literature on how to attack or manipulate AI (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). Recently, evidence appeared that such attacks occur on real-world systems (Grosse et al., 2024b; McGregor, 2021; ARVA, ). Such reports include, but are not limited to, malicious AI dependencies [^1] https://avidml.org/database/, 2023-V015, bypassed malware detectors [^2] https://avidml.org/database/, 2023-V004, reverse engineering of existing models (Grosse et al., 2024b), and search engine poisoning [^3] https://incidentdatabase.ai/cite/88. Research confirms (Carlini et al., 2024a) such threats for widely deployed AI systems like large language models (LLMs) (Baier, 2024), and text-based attacks can be mined from the internet and thus manifest there as well (Shen et al., 2024). Yet, AI security incidents, as of now, seem not to occur widely (Grosse et al., 2024b). It would thus be an ideal moment to start collecting AI security incidents for a proactive, not reactive, approach.
Intuitively, an attacker warrants different information to be collected than other incidents. Examples for incident collection from non-AI security include the Common Vulnerabilities and Exposures (CVE) program (MITRE, d; Baker et al., 1999). CVE is an established industry standard, and similar to the aviation case above, the collected data enable analysis for deeper understanding (Kumar & Sharma, 2022), like for example trends in of frequencies of exploits. However, existing security taxonomies do not apply well to AI: In contrast to shareable and protected software, AI model specifics (such as architecture, learned weights, and hyperparameters) may affect vulnerabilities (Biggio & Roli, 2018; Gu et al., 2019). Model specifics may also be the intellectual property or business asset of the affected company (Grosse et al., 2023). Furthermore, taxonomies like the CVE program catalog vulnerabilities found in the wild. In contrast, AI model vulnerabilities are often well-known with unknown practical exploit status. Understanding the exact exploit would be more fruitful. AI security incident reporting should thus require more context for understanding AI-specific vulnerabilities and exploits not covered by conventional incident reporting. We thus need a broad consensus on the necessary content, scope and granularity of AI security incident reporting that has not yet been reached from scientific, industry and market surveillance perspective.
@@
Our position: A reporting taxonomy for AI security incidents is necessary and (1) should contain specific information on the AI system, the incident and its implications. (2) Its structure should be aligned with existing reporting approaches and regulatory disclosure obligations. (3) The content should be of sufficient granularity for industrial practitioners, researchers and market surveillance authorities to derive required action or, in some cases, even reproduce the incident. The taxonomy should further be easy to use for reporters, and contain machine readable, properties that however depend on the exact implementation. Finally, we present this draft to spark further discussions in the community.
Outline. We first define relevant terminology based on existing standards or works, and review related concepts and approaches (Section 2). Afterwards, we describe how we derived our taxonomy by unifying related work and expert opinions with iterative rounds (Section 3). We present the resulting scheme in Section 4, and review limitations, implementation, implications and future work (Section 5).
Before concluding with the impact statement, we summarize alternative views (Section 6).
@@
Related Work. To the best of our knowledge, there is no established procedure for AI security incident reporting. Cattell & Ghosh (2024) propose a CVE-like framework for adjudicating AI flaws more generally (e.g., incident reporting for undesired behaviors of AI systems which is not focused on security) as well as other improvements (such as extended model cards). In this sense, Cattell & Ghosh (2024) is orthogonal to our work. In addition, Fazelnia et al. (2024) proposed and MITRE announced a general AI incident sharing initiative [^4] https://ai-incidents.mitre.org concurrently to us. While both proposals bear similarities to ours, we suggest a more detailed treatment of the intended purpose, supply chains, and existing security measures. However, Fazelnia et al. (2024) focus on a minimum set of requirements and also propose an AI bill of materials (AIBOM). As AIBOMs summarize all components of AI models (Fazelnia et al., 2024; Bennet et al., 2024), our taxonomy would also benefit from an AIBOM standard.
@@ 2 Background
We first motivate our work based on the legal background. Afterwards, we introduce, based on existing standards, relevant definitions. We then review existing works describing AI systems, in AI security, before we discuss existing infrastructure to report AI incidents and security vulnerabilities.
2.1 Incident reporting as regulatory requirement
From a legal perspective, there are security-related requirements applicable under Article 55 (1c) of the European AI Act [^5] https://eur-lex.europa.eu/legal-content/EN/TXT/PDF/?uri=OJ:L_202401689. More specifically, commitment 17 on serious incident reporting of the general-purpose AI code of practice, currently under development [^6] https://digital-strategy.ec.europa.eu/en/library/second-draft-general-purpose-ai-code-practice-Dpublished-written-independent-experts, has a direct relation to our work: According to these documents, providers of general-purpose AI systems will be obliged to report relevant incident information such as affected models, root causes, or fail-safe measures in case serious incidents occur. With our taxonomy, we aim to cover security-related serious incidents. Our goal is to further to sharpen content, scope, and granularity of future reporting requirements. Especially as market surveillance authorities such as the European AI Office will have to issue further guidance on their operationalization, we want to initiate and inform a practical debate.
@@ 2.2 Definitions
Consequently, we review core terminology based on existing standards and use them to construct a definition of “AI security incident.” The complete list of all required definitions is in Appendix A. Here, we only revisit the three most important definitions: firstly an “AI system”, secondly “security”, and thirdly an “AI security incident” (the target of our collection). We first start with the definition of the “AI system”, derived from the OECD (a) definition.
@@
AI system. An AI system is a machine-based system that, for explicit or implicit objectives, infers, from the input it receives, how to generate outputs such as predictions, content, recommendations, or decisions that can influence physical or virtual environments. Different AI systems vary in their levels of autonomy and adaptiveness after deployment.
We look to this definition of “AI system” as a foundation for our work for two main reasons. Firstly, the relevant OECD recommendation also identifies that “AI systems” and the “AI models” they embed are distinct. This allows a potential AI security incident report to identify if an exploit is scoped strictly to the “mathematical construct” that is a model, or interacts with the larger AI system. Secondly, the definition is used by or influences those of other intergovernmental organizations (such as the United Nations) and prominent market regulations relevant to incident reporting (such as the EU AI Act) (OECD, a).
We also recognize “security” as broadly defined in ISO/IEC 27000 by ISO (2018).
@@
We further additionally review the definitions of confidentiality, integrity, and availability, also known as the “CIA triad”. This well established concept is also used in scientific work covering AI security (Biggio & Roli, 2018).
@@
Importantly, this definition of “security” based primarily on the CIA triad is shared by other ISO/IEC standards specific to AI. [^7] For example: ISO/IEC 42001 Information technology — Artificial intelligence — Management system; ISO/IEC TR 24028 Information technology — Artificial intelligence — Overview of trustworthiness in artificial intelligence; ISO/IEC TR 24368 Information technology — Artificial intelligence — Overview of ethical and societal concerns
Based on these foundational terms, we define an “AI security incident”, extending the existing definition of “AI incident” by the OECD (2024) and considering the above CIA security triad aspects shared across ISO/IEC standards.
@@
We use these definitions to avoid, for example, confusion between general security and AI-specific security incidents (Bieringer et al., 2022). Also note that the proposed definition of an AI security incident includes potential compromises, hence vulnerabilities or red teaming results. It may be desirable to formulate this more strictly later-on.
@@ 2.3 Describing AI systems
Since the reporting of AI security incidents should be consistent with the overall documentation of an AI system, our proposed taxonomy builds on and expands prior work on model cards. After their initial proposal by Mitchell et al. (2019), these have been widely adopted by model providers and are used by various stakeholder groups to access information about a model and conclude corresponding action. Derivative templates have been proposed as fact sheets (Arnold et al. (2019)), datasheets for datasets (Gebru et al. (2021)), method cards (Adkins et al. (2022)), risk cards (Derczynski et al. (2023)), use case cards (Hupont et al. (2024)), and AI cards (Golpayegani et al. (2024)). While most of these approaches require the documentation of general metadata about the model (e.g., model type or model version), their content related to security is often limited and only covers implemented security measures, if at all. Yet, they provide suitable guidance on how to describe the underlying system affected by the incident. This is why the descriptions of AI systems affected by incidents within our proposed taxonomy (Section 4.2) are based on this body of literature.
@@ 2.4 AI security
Similar to any other (digital) system, an AI system can be targeted by a plethora of attacks (Biggio & Roli, 2018; Cinà et al., 2023; Gu et al., 2019; Lei et al., 2023; Oliynyk et al., 2023; Shen et al., 2024; Shumailov et al., 2021; Yao et al., 2024). To understand how AI is vulnerable, in the following, we give a very high-level overview based on the previously discussed CIA triad defined in the previous Section 2.2.
Poisoning attacks alter the training data, samples, or labels to decrease the overall performance of the classifier (Barreno et al., 2006; Cinà et al., 2023; Zhang et al., 2020). They, thus, target the integrity via the performance, or availability of the AI model if performance becomes too low.
Table 1: Existing databases and collections of AI (security) incidents. We list the kind of support party (Government, Industry, Non-Profit, and Academia) and incident numbers as of June 14th, 2024. We list the number of entries in each database and whether incidents cover the topics of ethics, security, AI security, or offensive use of AI within security. We then review which databases contain reproducible incidents. We then revisit the sources of the incidents: research papers, media, anonymous reports, or platforms where registered users can submit reports. We finally report whether the databases are machine-readable and linked with the CVE databases. The AVID database is linked to NVD, MITRE, and AIID.
@@
MITRE case studies (MITRE, a) G 22 JSON Python CVE
AIRD (MITRE, ) I 4 CSV Python CVE
Model stealing copies the model without the owner’s consent (Carlini et al., 2024b; Lowd & Meek, 2005; Oliynyk et al., 2023; Tramèr et al., 2016), targeting confidentiality.
@@
More recent attacks include backdoors (Cinà et al., 2023; Gu et al., 2017; Yang et al., 2024; Yu et al., 2022), weight perturbations (Breier et al., 2018; Neti et al., 1992), adversarial initialization (Grosse et al., 2020; Liu et al., 2019), and sponge attacks (Shumailov et al., 2020). The exact nature and scope of describing and tagging a specific attack in a practical context is a research question itself and thus left for future work. There are, however, established frameworks like the MITRE ATLAS [^8] https://atlas.mitre.org/matrices/ATLAS addressing this challenge in practice. An alternative route is to use existing research frameworks, for example by Biggio & Roli (2018).
2.5 AI security incident reporting frameworks
@@
Although there are reports of real-world AI security incidents as discussed in the introduction (Grosse et al., 2024b; McGregor, 2021; ARVA, ), there are not many security incident databases for AI systems. We summarize related databases in Table 1. Most databases cover not only AI security incidents but also other types of incidents, from unethical applications of AI to security; some include offensive use of AI (OECD, b; McGregor, 2021; ARVA, ; LVEs, ; MITRE, a). To enable conclusions specific to AI security, databases should be more focused. This is the case for fewer databases, which are narrow but however collect non-AI security (Huntr, ) or offensive AI usage (Walker et al., 2024). To the best of our knowledge only Grosse et al. (2024b) focus on AI security, with only a small sample of 32 incidents. In addition, most databases rely on publicly available information, like research papers and media reports (OECD, b; McGregor, 2021; ARVA, ; LVEs, ; MITRE, a). To enable the collection of AI security-relevant information, reports should be used that share relevant details. Yet, few databases rely on direct submissions (Huntr, ; LVEs, ; MITRE, ), including, for example, bug bounties. Even fewer use anonymous reports (Grosse et al., 2024b). Many collections also fall short in the underlying taxonomy that would capture relevant information for systematic analysis benefiting all stakeholders. While there are approaches with taxonomies, like the AI Incident Database (AIID) (McGregor, 2021), their focus should ideally be more narrow to allow security-specific conclusions. A suitable approach is the concurrently developed AI incident sharing initiative by MITRE [^9] https://ai-incidents.mitre.org. In contrast to this initiative, we advertise sharing mode details for analysis.
@@
Still, many databases provide guidelines on collected information and emphasize other relevant features. These include machine readability (seven of nine databases allow direct export of data), links to other vulnerability databases like CVE (four of nine), and lastly direct integration via for example APIs with programming languages for data analysis (five of nine). The latter is especially relevant to research, where corresponding frameworks in non-AI security have enabled more detailed analysis of the threat surface (Neuhaus & Zimmermann, 2010; Li et al., 2017).
2.6 Security incident reporting frameworks
Frameworks for collecting, classifying, and reporting IT security incidents began to take shape in 1999 with the seminal work on the Common Vulnerability and Exposures (CVE) database (Baker et al., 1999). The CVE Program (CVE, ) has since matured and achieved widespread adoption. Building on the success of CVE, complementary initiatives evolved.
Common Vulnerability Scoring System (CVSS) (NIST, a; Scarfone & Mell, 2009) was developed to assess the severity of vulnerabilities. As a result, the National Vulnerability Database (NVD) (NIST, b) now contains CVSS scores for all published CVE records.
Common Weakness Enumeration (CWE) (MITRE, b) lists software and hardware weaknesses that could potentially evolve into vulnerabilities. There may be a connection between CVE and CWE records, as a reported CVE vulnerability could be the result of a documented CWE weakness.
MITRE ATT&CK (MITRE, b, c) is a knowledge base of attack tactics and techniques. It relates to CVE and CWE in that ATT&CK entries may exploit reported vulnerabilities and weaknesses.
Previous research has highlighted that security vulnerabilities in general and those specific to AI are fundamentally different (NIST, 2023; Cattell & Ghosh, 2024). We agree with this position, yet acknowledge that the CVE project can also be relevant for AI models. For example, CVE-2023-25661101010https://nvd.nist.gov/vuln/detail/CVE-2023-25661 reports a Tensorflow-related vulnerability that could have implications for AI models. However, the scope of CVE, and its associated taxonomy, differs from our approach. The CVE project primarily focuses on IT-related vulnerabilities, while our AI security incident taxonomy is specifically designed to address incidents related to AI models.
Our work aims to translate subject matter expert input into a comprehensive, self-contained AI security incident taxonomy. This approach mirrors the CVE model, where reported vulnerabilities are reviewed and eventually added to the CVE database. This also means that our project can benefit from lessons learned throughout the evolution of the CVE project. For instance, as the volume of AI security incidents is expected to grow steadily, a labeling schema is required to uniquely identify incidents that are classified according to our proposed taxonomy. Additionally, the process of reviewing and labeling incidents could become resource-intensive. Therefore, a governance body — similar to the CVE Numbering Authorities (CNA) (MITRE, a) — must be established to oversee this task.
If our proposed taxonomy gains acceptance within both the AI and security communities, further initiatives can be explored such as integrating the CVE and our taxonomy, assigning severity scores similar to CVSS, or abstracting AI security incident instances into more generalized weaknesses, similar to those defined by CWE.